# Importance-Aware Robust Semantic Transmission for LEO Satellite-Ground Communication

### v1

@@ VI-A Dataset and Simulation Parameters
VI-B Comparison Schemes
VI-B1 Deep JSCC
Deep JSCC [33] adopts an end-to-end deep neural network framework that directly maps image pixels to channel input symbols, effectively bypassing the conventional processes of image compression and channel coding. Typically structured as an autoencoder, it comprises a CNN-based encoder and decoder, with a non-trainable channel simulation module—commonly modeled as an Additive White Gaussian Noise (AWGN) or Rayleigh fading channel—embedded between them.
VI-B2 WITT
@@ VI-B3 SS-DDPM
VI-C Performance Metrics
VI-C1 PSNR
Peak signal-to-noise ratio (PSNR) is an index used to measure the quality of images or signals. PSNR is mainly calculated based on the mean square error at the pixel level. Because of its simplicity and objectivity, it is widely used in image compression, denoising, and quality monitoring during transmission [18, 33, 34]. The higher the PSNR value, the higher the similarity between the two images, the better the quality. It is calculated as
@@ VI-C1 PSNR
VI-C2 SSIM
Structural Similarity Index Measure (SSIM) is a perceptual metric developed to approximate the characteristics of human visual perception when assessing the quality of digital images. Unlike traditional pixel-based methods, SSIM evaluates the similarity between two images based on three key components: luminance, contrast, and structural information. This comprehensive approach enables SSIM to more accurately reflect perceived visual quality, particularly in scenarios where structural fidelity is crucial, such as image enhancement, super-resolution reconstruction, and facial image processing [34, 43, 44]. The SSIM index ranges from 0 to 1, with values closer to 1 indicating a higher degree of similarity between the compared images. It is calculated as
@@ VI-C2 SSIM
VI-D Results Analysis
VI-D1 Module Performance Verification
@@
Table IV presents a comparative analysis of the inclusion of the SME module. Although its integration does not result in notable changes in overall accuracy or general performance metrics, the presence of SME substantially enhances the quality of segmentation in edge regions. This improvement is attributed to the module’s ability to refine boundary information handling in comparison to the original scheme, thereby exerting a significant positive influence on the effectiveness of segmentation.
@@
Figure 5 presents a performance comparison of remote sensing image transmission under conditions with and without the application of the proposed fuzzy filtering scheme. This scheme is designed to optimize bandwidth usage while preserving the visual integrity of task-critical regions. In Figure 5(a), the transmitted image size is 160 KB, yielding an overall PSNR of 38.648 dB and a task-region PSNR of 39.077 dB. Although the image quality remains uniformly high, the substantial data size results in elevated transmission costs. In contrast, Figure 5(b) demonstrates the effectiveness of fuzzy filtering, reducing the image size to 117 KB—a compression rate of approximately 26.9 %. While the overall SNR decreases to 32.747 dB and background details exhibit visible degradation, the task-region PSNR remains stable at 39.067 dB, closely matching the unfiltered result. This outcome suggests that the scheme maintains the perceptual quality of the task-critical area while significantly improving transmission efficiency. The design encapsulates the core principle of task-oriented semantic communication: selectively suppressing redundant visual information to ensure semantic goal completion under constrained bandwidth conditions.
@@
Figure 6: Performance comparison of whether to adopt the stacking scheme.
VI-D2 Performance Comparison
Figure 7 presents a comparative evaluation of image quality metrics across various SNR conditions for multiple image transmission schemes. Specifically, Figure 7(a) illustrates the PSNR performance trends. PSNR values increase with rising SNR levels for all methods, with the proposed IRST scheme demonstrating consistent superiority across the entire SNR spectrum. Notably, it achieves nearly 30 dB under low SNR conditions, underscoring its resilience in noisy environments. Figure 7(b) depicts SSIM variation across SNR levels. The IRST scheme, employing a stacked architecture, maintains high SSIM values from low to high SNR, indicating enhanced structural fidelity. Moreover, its SSIM values exhibit minimal fluctuation throughout the SNR range, reflecting robust adaptability. This stability arises from the dynamic participation of constituent models: under low SNR, more models are engaged during channel encoding and decoding to improve robustness; under high SNR, fewer models are activated, thereby reducing computational complexity and latency.
@@ VI-D3 Adaptation Comparison
Overall, the experimental findings affirm that the IRST architecture possesses robust generalization across datasets. It effectively sustains image quality under diverse channel conditions and data discrepancies, highlighting its suitability for real-world deployment in remote sensing communication. By contrast, Deep JSCC’s limited capacity to extract heterogeneous semantic features contributes to its poor adaptability when tested outside its training domain.
@@ VII Conclusion
The integration of SemCom with satellite networks offers significant potential for future intelligent communication systems. In this paper, we propose the IRST model in the scenario of limited bandwidth and dynamic channel changes. In this model, an SME scheme is first proposed to improve the accuracy of semantic segmentation. Then, based on the task requirements and the current channel conditions, we propose a semantic selection scheme to select the transmission content. Then, we design a stack-based SNR channel codec, which can perform dynamic channel coding according to the current SNR. Finally, we compare the proposed model with other schemes and the results show that the IRST model performs well under various conditions.

### v2

@@ VI-B Dataset and Simulation Parameters
VI-C Comparison Schemes
VI-C1 Deep JSCC
Deep JSCC [33] adopts an end-to-end deep neural network framework that directly maps image pixels to channel input symbols, effectively bypassing the conventional processes of image compression and channel coding. Typically structured as an autoencoder, it comprises a Convolutional Neural Network (CNN)-based encoder and decoder, with a non-trainable channel simulation module—commonly modeled as an Additive White Gaussian Noise (AWGN) or Rayleigh fading channel—embedded between them.
VI-C2 WITT
@@ VI-C3 SS-DDPM
VI-D Performance Metrics
VI-D1 PSNR
Peak Signal-to-Noise Ratio (PSNR) is an indicator used to measure the quality of an image or signal. By comparing the pixel difference between the original image and the transmitted image, the distortion degree of the image is quantified. PSNR is mainly calculated based on the mean square error at the pixel level. Because of its simplicity and objectivity, it is widely used in image compression, denoising, and quality monitoring during transmission [19, 33, 34]. The higher the PSNR value, the higher the similarity between the two images, the better the quality. It is calculated as
@@ VI-D1 PSNR
VI-D2 SSIM
Structural Similarity Index Measure (SSIM) is a perceptual metric developed to approximate the characteristics of human visual perception when assessing the quality of digital images. Unlike traditional pixel-based methods, SSIM evaluates image similarity from three aspects : brightness, contrast and structure, which is closer to the human visual system. In the process of image transmission, SSIM can more accurately reflect whether the image looks the same, even if the pixels are slightly different. This comprehensive approach enables SSIM to more accurately reflect perceived visual quality, particularly in scenarios where structural fidelity is crucial, such as image enhancement, super-resolution reconstruction, and facial image processing [34, 46, 47]. The SSIM index ranges from 0 to 1, with values closer to 1 indicating a higher degree of similarity between the compared images. It is calculated as
@@ VI-D2 SSIM
VI-D3 LPIPS
Learned Perceptual Image Patch Similarity (LPIPS) is a perceptual metric designed to evaluate the visual similarity between two images from the perspective of human perception. Unlike traditional pixel-wise metrics such as PSNR or SSIM, LPIPS compares deep feature representations extracted from pretrained neural networks, thereby capturing high-level semantic differences that are more aligned with human judgment. Due to its ability to reflect perceptual quality rather than mere pixel fidelity, LPIPS has been widely adopted in tasks such as image generation, super-resolution, style transfer, and adversarial robustness evaluation [48, 49]. A lower LPIPS score indicates higher perceptual similarity between the reference and the distorted image. The LPIPS metric is formally calculated as
$LPIPS(x,x_{0})=\sum_{l}w_{l}\cdot\|f_{l}(x)-f_{l}(x_{0})\|_{2}^{2},$ (30)
where $f_{l}(\cdot)]$denotes the feature map extracted from layer $l$ of a pre-trained network, and $w_{l}$ represents the learned weight for each layer. $x$ and $x_{0}$ are the input and reference images, respectively.
VI-E Results Analysis
VI-E1 Module Performance Verification
@@
Table IV presents the comparative analysis of whether to use SME module, including accuracy, PSNR, SSIM and boundary Intersection over Union (IoU). The results indicate that although SME does not significantly improve overall accuracy, it significantly improves the quality of boundary segmentation. This is because the encoder part of SegNet uses a multi-layer maximum pooling operation. Although this enhances the translation invariance and robustness of the model, it also significantly reduces the spatial resolution of the feature map, making the boundary details blurred or lost during the downsampling process, resulting in a decrease in recognition rate. The SME improves the accuracy of location recognition by supplementing the points of the surrounding environment.
@@
Boundary IoU 0.718 0.638
Figure 5 presents a performance comparison of remote sensing image transmission under conditions with and without the application of the proposed fuzzy filtering scheme. This scheme is designed to optimize bandwidth usage while preserving the visual integrity of task-critical regions. In Figure 5(a), the transmitted image size is 160 KB, yielding an overall PSNR of 38.648 dB and a task-region PSNR of 39.077 dB. Although the image quality remains uniformly high, the substantial data size results in elevated transmission costs. In contrast, Figure 5(b) demonstrates the effectiveness of fuzzy filtering, reducing the image size to 117 KB—a compression rate of approximately 26.9 %. While the overall SNR decreases to 32.747 dB and background details exhibit visible degradation, the task-region PSNR remains stable at 39.067 dB, closely matching the unfiltered result. This outcome suggests that the scheme maintains the perceptual quality of the task-critical area while significantly improving transmission efficiency. The design encapsulates the core principle of task-oriented SemCom: selectively suppressing redundant visual information to ensure semantic goal completion under constrained bandwidth conditions.
@@
Figure 6: PSNR and SSIM comparison under varying SNR conditions between IRST with stacking scheme and baseline without stacking optimization.
VI-E2 Performance Comparison
Figure 7 presents a comparative evaluation of image quality metrics across various SNR conditions for multiple image transmission schemes. Specifically, Figure 7(a) illustrates the PSNR performance trends. PSNR values increase with rising SNR levels for all methods, with the proposed IRST scheme demonstrating consistent superiority across the entire SNR spectrum. Notably, it achieves nearly 30 dB under low SNR conditions, underscoring its resilience in noisy environments.
Figure 7(b) depicts SSIM variation across SNR levels. The IRST scheme, employing a stacked architecture, maintains high SSIM values from low to high SNR, indicating enhanced structural fidelity. Moreover, its SSIM values exhibit minimal fluctuation throughout the SNR range, reflecting robust adaptability. This stability arises from the dynamic participation of constituent models: under low SNR, more models are engaged during channel encoding and decoding to improve robustness; under high SNR, fewer models are activated, thereby reducing computational complexity and latency.
Figure 7(c) further complements the evaluation by presenting LPIPS trends across varying SNR levels. The proposed IRST scheme consistently achieves the lowest LPIPS scores across the entire SNR range, highlighting its superior perceptual fidelity. Even under severely degraded conditions (e.g., SNR = - 10 dB), IRST maintains a perceptual advantage over competing methods such as SS-DDPM, Deep JSCC, and WITT. This performance underscores IRST’s ability to preserve visually meaningful features despite channel noise, further validating the effectiveness of its adaptive model stacking strategy.
@@ VI-E3 Adaptation Comparison
Figure 8(c) presents the LPIPS performance of the four image transmission schemes under varying SNR conditions. The proposed IRST method consistently achieves the lowest LPIPS values throughout the SNR range, indicating superior perceptual similarity and robustness to noise. WITT is second, particularly in the medium to high SNR regions, where it maintains competitive perceptual quality. SS-DDPM demonstrates moderate performance, showing effective denoising but limited perceptual fidelity in low SNR scenarios. Meanwhile, Deep JSCC also exhibits the high LPIPS scores, especially under low SNR conditions, reflecting its limited ability to preserve semantic and perceptual features due to its simplistic architecture.
Overall, the experimental findings affirm that the IRST architecture possesses robust generalization across datasets. It effectively sustains image quality under diverse channel conditions and data discrepancies, highlighting its suitability for real-world deployment in remote sensing communication. In contrast, the limited capacity of Deep JSCC to extract heterogeneous semantic features contributes to its poor adaptability when tested outside of its training domain.
@@
(c) LPIPS performance comparison in using different dataset.
@@ VI-E4 Visualization Result
VI-E5 Complexity Comparison
Table V presents a comparative analysis of model complexity across different architectures. The first row reports the memory footprint associated with model parameters, while the second row details the number of Multiply-Accumulate Operations (MACs), and the third line reports the GPU running time. As illustrated, both Deep JSCC and SS-DDPM primarily rely on CNN, leveraging parameter sharing and local receptive fields. This design results in relatively low memory consumption and computational cost. However, such architectures tend to struggle with preserving resolution when processing high-definition images under low SNR conditions. Furthermore, since SS-DDPM is based on a diffusion model for image reconstruction, its runtime is significantly longer than other methods.
In contrast, IRST and WITT utilize the Swin Transformer architecture, which incorporates hierarchical feature extraction and cross-regional modeling. These mechanisms significantly increase both memory usage and computational complexity, yet yield superior generative performance. Notably, IRST employs a dynamic SNR-aware channel codec, enabling more efficient memory utilization compared to WITT, which uses separately designed codecs for different channel conditions. Furthermore, because IRST skips redundant coding paths under high SNR conditions, it significantly reduces runtime, making it suitable for resource-constrained scenarios.
Overall, the proposed IRST model achieves a balanced trade-off between complexity and output quality, maintaining computational stability while adapting effectively to varying channel environments.
TABLE V: Complexity Comparison
Model IRST WITT Deep JSCC $S$-DDPM
Parameter Size (M) 19.89 21.72 1.56 5.11
FLOPs (GMac) 32.45 32.82 2.62 6.73
GPU Runtime (s) 4.97 6.79 5.51 8.04
@@ VII Conclusion
The integration of SemCom with satellite networks offers significant potential for future intelligent communication systems. In this paper, we propose the IRST model in the scenario of limited bandwidth and dynamic channel changes. In this model, an SME scheme is first proposed to improve the accuracy of semantic segmentation. Then, based on the task requirements and the current channel conditions, we propose a semantic selection scheme to select the transmission content. Meanwhile, we design a stack-based SNR channel codec, which can perform dynamic channel coding according to the current SNR. Finally, we compare the proposed model with other schemes and the results show that the IRST model performs well under various conditions.
Although the IRST model proposed in this paper demonstrates significant improvements in the robustness and efficiency of satellite-ground SemCom, several limitations remain that warrant further investigation. First, the current validation is performed primarily on remote sensing image tasks, and the model adaptability to other modalities (such as speech, text, and video) has not been thoroughly explored, thus constraining its generalization capability. Moreover, channel modeling and feedback mechanisms are based on idealized assumptions, without fully accounting for the dynamic variations and latency inherent in real-world satellite-ground links, including Doppler effects and energy consumption, which may compromise the accuracy of semantic selection. Future research directions include extending the IRST framework to support multimodal data in order to enhance system versatility; incorporating reinforcement learning and adaptive strategies to enable end-to-end semantic optimization and dynamic policy adjustment; and conducting deployment verification in real world or high-fidelity simulated satellite-ground environments to rigorously assess the stability and practical applicability of the model under complex conditions.